OpenStudy (anonymous):

Which one of the equations has a graph with no x-intercepts? A) y = -x2 B) y = x2 - 1 C) y = -x2 - 1 D) y = -x2 + 1

OpenStudy (anonymous):

It would be the graph that has no solution when y=0. C looks about right, since \[0\le x^2 \le \infty \]And the negative sign means -x^2 will always be negative, and -1 off that would mean its just a lower negative number. It would never touch 0.